OpenStudy (anonymous):

Solve the equation on the interval 0,2pi 1-cos=1/2

OpenStudy (saifoo.khan):

-cosx = 1/2 -1 -cosx = -1/2 cos x = 1/2

OpenStudy (saifoo.khan):

Can you solve from here? @axlopez23

OpenStudy (anonymous):

write your question again............

OpenStudy (anonymous):

is it 1-cosx=1/2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

pi/3, 5pi/3?

OpenStudy (anonymous):

cos(x)=1/2 cos(x)=cos((pi)/3)or cos(5pi/3)

OpenStudy (anonymous):

ya your answer is right

OpenStudy (anonymous):

ok how about sin(3x)=-1? sin is -1 at 3pi/2

OpenStudy (maheshmeghwal9):

so x = pi/2

OpenStudy (anonymous):

so then would I use the +pi k formula?

OpenStudy (maheshmeghwal9):

hmm..................?

OpenStudy (anonymous):

sin(3x)=-sin(pi/2) sin3x=sin(pi+pi/2) or sin(2pi-pi/2) sin3x=sin(3pi/2)

OpenStudy (anonymous):

sorry

OpenStudy (maheshmeghwal9):

so x = pi/2 @nitz ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

hmm ok so I know pi/2 is where sin = -1, that would be my only solution in the interval right?

OpenStudy (anonymous):

i meant 3pi/2. Would pi/2 be part of the solution as well?

OpenStudy (maheshmeghwal9):

no only pi/2 is ur answer becoz u had to find the value of x which is pi/2 which again will the point at which sin 3x will become -1 & the angle at which the value of above function is -1 is 3x=3pi/2. so it is the angle nt ur answer ur answer is x=pi/2

OpenStudy (anonymous):

ok thanks. im stuck on this last one of this type as well, tan x/2= root3. I came up with x=1, x=-1/2.